OpenStudy (anonymous):

Solve graphically y = 2x - 6 and y = -2x + 10. Write down the co-ordinates of the point of intersection of the two lines

OpenStudy (amistre64):

this site really isnt designed to draw a graph on ..

OpenStudy (anonymous):

i really just need help finding the coordinates so i can graph it

OpenStudy (amistre64):

im curious as to whether the y component is centered between you intercepts :)

OpenStudy (anonymous):

set the lines equal and solve \[2x-6=-2x+10\]

OpenStudy (amistre64):

plug a point at -6 and slope up 2 right 1 and connect the dots

OpenStudy (anonymous):

\[4x=16\] \[x=4\] and since \[x=4\] then \[y=2\times 4-6=2\]

OpenStudy (amistre64):

y should be 2 if i see it right :)

OpenStudy (amistre64):

yep ... i sawed it right lol

OpenStudy (anonymous):

lol sweet thanks

OpenStudy (anonymous):

to find the points of intersection, set the equations equal to each other. \[2x -6 = -2x + 10\] \[4x = 16\] x= 4 Then plug in the x coordinate to one of the equations. y= 2(4) -6 = 2 The point is ( 4. 2)

OpenStudy (anonymous):

wolfram is nice, but to solve \[2x-6=-2x+10\] rather unnecessary